A +1C charge Is placed at position X=-2m,  while a - 3C charge is placed at position X=3m. A)Find net electric field at the origin.B) Find the net force (magnitude and direction) on a 1nC test particle placed at the origin. Please write neatly!
Answer #1

Dear student,

Find this solution, and RATE IT ,If you find it is helpful .your rating is very important to me.If any incorrectness ,kindly let me know I will rectify them soon.

Thanks for asking ..q=+1C, @ (-2m, o) 9=-30 @ 3m, 9) +4 Er -12 2m 3m + N/ a) Enet = E14 E2= K + K221 = 9 X 10 and 22 (2) = Enet = 5.25x19 alwe to